OpenStudy (anonymous):

D=1/4(e-3f) solve for e

OpenStudy (anonymous):

D = 1/[4*(e-3f)] 4*(e-3f) = 1/D 4e-12f = 1/D 4e = (1/D) + 12f e= [1/(4D)] + 3f i hope this is what your looking.